# Intention-Guided Cognitive Reasoning for Egocentric Long-Term Action Anticipation

### v1

@@ Introduction
The first category emphasizes enhancing visual modeling through improved feature extraction, aggregation strategies, and advanced spatiotemporal architectures. For instance, lightweight models (Hussein, Gavves, and Smeulders 2019a, b; Das and Ryoo 2022; Nawhal, Jyothi, and Mori 2022; Ishibashi et al. 2023) have been proposed to boost efficiency and representation quality. Additionally, Zhang et al. (Zhang et al. 2024a) introduced object-centric feature extraction using natural language queries, combined with dual-Transformer modules for encoding and temporal aggregation. Despite these innovations, such methods typically suffer from limited capacity and an absence of external priors, resulting in poor generalization in complex, dynamic, and cluttered egocentric scenarios.
@@
Traditional LTA approaches apply visual encoders (e.g., CLIP (Radford et al. 2021), EgoVLP (Lin et al. 2022), EgoVideo-V (Pei et al. 2024)) directly to entire frames. To improve the capture of relevant visual details in egocentric video, we introduce an HOI-focused feature extraction strategy. Specifically, given a video segment $S_{k}\in\mathcal{V}_{\mathrm{obs}}$, we follow AntGPT (Zhao et al. 2023) and uniformly sample 444 frames per segment, denoted as $F_{k,T}$ with $T\in\{1,\ldots,4\}$. Each frame $F_{k,T}$ undergoes HOI region detection using the pretrained 100DOH detector (Shan et al. 2020), refined with high-resolution masks from SAM2 (Ravi et al. 2024), producing precise masks $R_{k,T}$ representing active HOI regions.
@@
4) Soft Overlong Punishment: A soft penalty $R_{Soft}$ discourages overly verbose outputs. If the generated length $L$ falls within $[L_{max}-L_{cache},L_{max}]$, the penalty decays linearly from 0 to 1-1- 1. If $L$ exceeds $L_{max}$, the penalty is fixed at 1-1- 1. Here $L_{max}$ denotes the maximum allowable length and $L_{cache}$ is the buffer length.
@@
where $\mathrm{ED}(\cdot,\cdot)$ denotes the edit distance function. The final accuracy reward is normalized to the range [0,1][0,1][ 0 , 1 ] as follows:
@@
[…]
@@
1) Ego4D (Grauman et al. 2022) is among the most comprehensive egocentric datasets available, capturing diverse daily-life scenarios. The v1 release includes 3,6703,6703 , 670 hours of footage, with the Forecasting subset containing 1,7231,7231 , 723 annotated clips (116116116 hours) labeled with 115115115 verbs and 478478478 nouns. The updated v2 dataset expands this subset to 3,4723,4723 , 472 clips (243243243 hours), annotated with 117117117 verbs and 521521521 nouns. We employed the official data splits from the v2 version for our experiments.
2) EPIC-Kitchens-55 (Damen et al. 2018) comprises egocentric cooking activities recorded in realistic kitchen settings. This dataset includes 555555 hours of densely annotated videos with 125125125 verbs and 352352352 nouns. Our experiments utilized standard data splits defined by EGO-TOPO (Nagarajan et al. 2020) to ensure comparability with previous studies.
3) EGTEA Gaze+ (Li, Liu, and Rehg 2018) consists of 868686 cooking-focused egocentric videos totaling 262626 hours. It provides annotations for 191919 verbs and 515151 nouns. Consistent with established evaluation protocols, we followed the official splits proposed in (Nagarajan et al. 2020).
@@
We adopted evaluation protocols specific to each dataset, ensuring consistency and comparability with prior works. For the Ego4D dataset, we adhered to the official evaluation approach, which specifies an observed context of 888 segments and requires predicting 555 distinct future action sequences, each comprising 202020 verb-noun pairs. We evaluated using edit distance (Damerau 1964) computed separately for verbs, nouns, and combined actions. The evaluation score is the minimal edit distance achieved between any of the predicted sequences and the ground-truth sequence.
@@
We employed a frozen EgoVideo-V encoder to extract visual embeddings from full frames and corresponding HOI regions. These dual-stream features were linearly fused and processed by a Transformer architecture consisting of 444 layers and 888 attention heads per layer to capture spatiotemporal dependencies. The feature extraction stage was trained using a batch size of 888, a learning rate of $8e-5$, and a plateau-based learning rate scheduler. For cognitive reasoning, we used Qwen2.5-VL-Instruct-7B (Bai et al. 2025) as the backbone, incorporating structured prompting (Figure 3) and trained end-to-end with Swift framework (Zhao et al. 2024) using our GRPO-based reinforcement learning objective. Experiments were conducted on six NVIDIA H20-SXM5-96GB GPUs.
Hyperparameters for fine-tuning included batch size of 242424, learning rate of $3e-6$, temperature of 0.90.90.9, KL divergence coefficient of 0.080.080.08, generation numbers of 555, buffer length of 256256256, and a maximum allowable length of 450450450. Reward function weights were set as follows: $\omega_{1}=0.90$, $\omega_{2}=0.10$, $\omega_{3}=0.85$, $\omega_{4}=0.05$, $\omega_{5}=0.05$, and $\omega_{6}=0.05$. The intention reward parameters were $\beta=0.8$ and $\gamma=40$. The full fine-tuning process consists of 500500500 steps, taking approximately 909090 GPU hours.
@@
Figure 4 presents learning dynamics under the adopted training configuration. (a) illustrates the progression of the overall reward, while (b) tracks the intention-specific reward; both panels show raw and smoothed curves to capture training trends. Training converges within 500500500 steps: rewards rise sharply during the initial 11001-1001 - 100 steps, followed by gradual improvement and stabilization between 100300100-300100 - 300 steps. Importantly, the intention reward closely mirrors the overall reward trajectory, suggesting that intention supervision is well aligned with the task objective. Although raw curves exhibit fluctuations typical of on-policy learning and exploration, smoothed trends indicate consistent performance gains without signs of instability. These results demonstrate the robustness of GRPO under our selected hyperparameters and validate the utility of explicit intention guidance in improving policy quality and convergence.
@@ Conclusion
Intention-Guided Cognitive Reasoning for Egocentric Long-Term Action Anticipation

### v2

@@ Introduction
The first category emphasizes enhancing visual modeling through improved feature extraction, aggregation strategies, and advanced spatiotemporal architectures. For instance, lightweight models (Hussein, Gavves, and Smeulders 2019; Das and Ryoo 2022; Nawhal, Jyothi, and Mori 2022; Ishibashi et al. 2023) have been proposed to boost efficiency and representation quality. Additionally, Zhang et al. (Zhang et al. 2024a) introduced object-centric feature extraction using natural language queries, combined with dual-Transformer modules for encoding and temporal aggregation. Despite these innovations, such methods typically suffer from limited capacity and an absence of external priors, resulting in poor generalization in complex, dynamic, and cluttered egocentric scenarios.
@@
Traditional LTA approaches apply visual encoders (e.g., CLIP (Radford et al. 2021), EgoVLP (Lin et al. 2022), EgoVideo-V (Pei et al. 2024)) directly to entire frames. To improve the capture of relevant visual details in egocentric video, we introduce an HOI-focused feature extraction strategy. Specifically, given a video segment $S_{k}\in\mathcal{V}_{\mathrm{obs}}$, we follow AntGPT (Zhao et al. 2023) and uniformly sample $4$ frames per segment, denoted as $F_{k,T}$ with $T\in\{1,\ldots,4\}$. Each frame $F_{k,T}$ undergoes HOI region detection using the pretrained 100DOH detector (Shan et al. 2020), refined with high-resolution masks from SAM2 (Ravi et al. 2024), producing precise masks $R_{k,T}$ representing active HOI regions.
@@
4) Soft Overlong Punishment: A soft penalty $R_{Soft}$ discourages overly verbose outputs. If the generated length $L$ falls within $L_{max}-L_{cache},L_{max}]$, the penalty decays linearly from 0 to 1-1. If $L$ exceeds $L_{max}$, the penalty is fixed at 1-1. Here $L_{max}$ denotes the maximum allowable length and $L_{cache}$ is the buffer length.
@@
where $\mathrm{ED}(\cdot,\cdot)$ denotes the edit distance function. The final accuracy reward is normalized to the range $[0,1]$ as follows:
@@
$S_{int}=\min\!\Bigl(\tfrac{1}{1+\exp[-\gamma(sim-\beta)]}\!\bigl/\!\tfrac{1}{1+\exp[-\gamma(1-\beta)]},\,1\Bigr),\quad$ (13)
@@
1) Ego4D (Grauman et al. 2022) is among the most comprehensive egocentric datasets available, capturing diverse daily-life scenarios. The v1 release includes $3,670$ hours of footage, with the Forecasting subset containing $1,723$ annotated clips ($116$ hours) labeled with $115$ verbs and $478$ nouns. The updated v2 dataset expands this subset to $3,472$ clips ($243$ hours), annotated with $117$ verbs and $521$ nouns. We employed the official data splits from the v2 version for our experiments.
2) EPIC-Kitchens-55 (Damen et al. 2018) comprises egocentric cooking activities recorded in realistic kitchen settings. This dataset includes $55$ hours of densely annotated videos with $125$ verbs and $352$ nouns. Our experiments utilized standard data splits defined by EGO-TOPO (Nagarajan et al. 2020) to ensure comparability with previous studies.
3) EGTEA Gaze+ (Li, Liu, and Rehg 2018) consists of $86$ cooking-focused egocentric videos totaling $26$ hours. It provides annotations for $19$ verbs and $51$ nouns. Consistent with established evaluation protocols, we followed the official splits proposed in (Nagarajan et al. 2020).
@@
We adopted evaluation protocols specific to each dataset, ensuring consistency and comparability with prior works. For the Ego4D dataset, we adhered to the official evaluation approach, which specifies an observed context of $8$ segments and requires predicting $5$ distinct future action sequences, each comprising $20$ verb-noun pairs. We evaluated using edit distance (Damerau 1964) computed separately for verbs, nouns, and combined actions. The evaluation score is the minimal edit distance achieved between any of the predicted sequences and the ground-truth sequence.
@@
We employed a frozen EgoVideo-V encoder to extract visual embeddings from full frames and corresponding HOI regions. These dual-stream features were linearly fused and processed by a Transformer architecture consisting of $4$ layers and $8$ attention heads per layer to capture spatiotemporal dependencies. The feature extraction stage was trained using a batch size of $8$, a learning rate of 8e58e-5, and a plateau-based learning rate scheduler. For cognitive reasoning, we used Qwen2.5-VL-Instruct-7B (Bai et al. 2025) as the backbone, incorporating structured prompting (Figure 3) and trained end-to-end with Swift framework (Zhao et al. 2024) using our GRPO-based reinforcement learning objective. Experiments were conducted on six NVIDIA H20-SXM5-96GB GPUs.
Hyperparameters for fine-tuning included batch size of $24$, learning rate of 3e63e-6, temperature of $0.9$, KL divergence coefficient of $0.08$, generation numbers of $5$, buffer length of $256$, and a maximum allowable length of $450$. Reward function weights were set as follows: $\omega_{1}=0.90$, $\omega_{2}=0.10$, $\omega_{3}=0.85$, $\omega_{4}=0.05$, $\omega_{5}=0.05$, and $\omega_{6}=0.05$. The intention reward parameters were $\beta=0.8$ and $\gamma=40$. The full fine-tuning process consists of $500$ steps, taking approximately $90$ GPU hours.
@@
Figure 5 presents learning dynamics under the adopted training configuration. (a) illustrates the progression of the overall reward, while (b) tracks the intention-specific reward; both panels show raw and smoothed curves to capture training trends. Training converges within $500$ steps: rewards rise sharply during the initial 11001-100 steps, followed by gradual improvement and stabilization between 100300100-300 steps. Importantly, the intention reward closely mirrors the overall reward trajectory, suggesting that intention supervision is well aligned with the task objective. Although raw curves exhibit fluctuations typical of on-policy learning and exploration, smoothed trends indicate consistent performance gains without signs of instability. These results demonstrate the robustness of GRPO under our selected hyperparameters and validate the utility of explicit intention guidance in improving policy quality and convergence.
@@ Conclusion
Acknowledgments
This work is supported by Shenzhen Science and Technology Program, No.:KQTD20240729102207002; Jiangsu Science and Technology Major Program, No.:BG2024041; the National Natural Science Foundation of China, No.:62376140 and No.:U23A20315; the Science and Technology Innovation Program for Distinguished Young Scholars of Shandong Province Higher Education Institutions, No.:2023KJ128; the Special Fund for Taishan Scholar Project of Shandong Province; the Major Key Project of Pengcheng Laboratory; the Fundamental Research Funds for the Central Universities, No.:HIT.DZJJ.2025048.
Intention-Guided Cognitive Reasoning for Egocentric
Long-Term Action Anticipation